### v2

# GNN 101: Visual Learning of Graph Neural Networks in Your Web Browser

Abstract.
Graph Neural Networks (GNNs) have achieved significant success across various applications. However, their complex structures and inner workings can be challenging for non-AI experts to understand. To address this issue, we present GNN101, an educational visualization tool for interactive learning of GNNs. GNN101 seamlessly integrates mathematical formulas with visualizations via multiple levels of abstraction, including a model overview, layer operations, and detailed animations for matrix calculations. Users can easily switch between two complementary views: a node-link view that offers an intuitive understanding of the graph data, and a matrix view that provides a space-efficient and comprehensive overview of all features and their transformations across layers. GNN101 not only demystifies GNN computations in an engaging and intuitive way but also effectively illustrates what a GNN learns about graph nodes at each layer. To ensure broad educational access, GNN101 is open-source and available directly in web browsers without requiring any installations.
Graph Neural Networks, Educational Visualization, Interactive Visualization, VIS4ML
ccs: Human-centered computing Visualizationccs: Human-centered computing Information visualization
Figure 1. (a) Traditional methods for learning Graph Neural Networks (GNNs) primarily rely on mathematical formulas, abstract diagrams, and programming code, which do not effectively convey the complete inner workings of GNNs. (b) GNN101 enhances understanding of GNNs through complementary views (B1), a hierarchical breakdown of details (B2), and the integration of mathematical concepts with visualizations (B3).
[…]
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
1. Introduction
Graph Neural Networks (GNNs) offer powerful capabilities for analyzing graph-structured data (e.g., social networks, molecular graphs). This type of data is often difficult to be effectively modeled by traditional machine learning models, which are mainly designed for non-graph data such as images and text. Therefore, GNNs have earned increasing popularity, especially in AI4Science research that leverages the inherent graph structures in scientific data to drive discoveries (e.g., new material design and drug development) (Huang et al., 2024; Wang et al., 2022).
However, learning GNNs can be more challenging than ML models for non-graph data, due to the unique characteristics of graphs and the complexities of graph-based computations. Unlike structured data such as images or tables, graphs are inherently non-Euclidean, meaning their data representation cannot be neatly organized into fixed-size grids. This irregularity makes it difficult to interpret standard computational operations like convolutions or pooling. Meanwhile, graphs can be sparse or have massive sizes, posing significant computational and memory challenges. To address these issues, techniques like graph sampling, mini-batching, and message passing have been developed, which introduce additional complexity for learning GNNs.
Recently, many educational resources of GNNs have been developed, covering formats ranging from online blogs to lecture videos and computational notebooks. Despite the great values provided by these educational resources, they often rely heavily on static illustrations (e.g., diagrams or mathematical equation) for explaining GNNs. Therefore, their focus remains on high-level overviews and theoretical frameworks, with limited emphasis on presenting the detailed inner workings of GNNs in a more intuitive and interactive manner.
Interactive visualization has long been recognized as an effective method for understanding complex algorithms and machine learning models (Wang et al., 2020; Hundhausen et al., 2002). As such, it offers a promising approach to facilitate more effective learning of GNNs by making their intricate processes and concepts more accessible and intuitive. However, visualizing GNNs can be challenging due to the inherent complexity of graph structures and the extensive mathematical details involved in GNN computations. While recent research has proposed visualization systems for GNNs, such as GNNLens (Jin et al., 2022) and CorGIE (Liu et al., 2022), these tools are primarily aimed at AI developers for debugging and improving models. Therefore, they focus on specific aspects related to GNN performances, such as error patterns (Jin et al., 2022) and embedding qualities (Liu et al., 2022), and provide limited support for beginners to learn about GNNs. Most relevant to this paper are the two interactive articles from Distill (Sanchez-Lengeling et al., 2021; Daigavane et al., 2021), which effectively use interactive visualizations to explain various core GNN concepts (e.g., adjacency matrix, convolution, pooling). However, the two papers primarily focus on individual concepts using abstract examples, and do not bridges theoretical understanding with practical application, such as the execution of actual GNN predictions.
To address these limitations and provide a more effective learning experience for GNNs, we introduce GNN101, the first interactive visualization tool for learning GNNs. The design of GNN101 is informed by a thorough analysis of existing GNN educational materials. To address the challenges in GNN visualization, GNN101 features a seamless linkage between mathematical formulas and visualizations via different levels of GNN details and provides two complementary views to enhance the learning experience. Specifically, it integrates a model overview, layer operations, and detailed animations for matrix calculations with smooth animations. It complements the advantages of a node-link view, which offers an intuitive representation of the graph structure, and a matrix view, which delivers a space-efficient and comprehensive overview of all features. We are currently deploying GNN101 in three courses that involve the teaching of GNNs at three different universities. This deployment has led to usage scenarios and observational studies that demonstrate the usability of GNN101 and generate design lessons for educational AI visualization tools.
As an open-sourced, web-based tool, GNN101 can be accessed without any installation, ensuring easy and widespread availability. It operates locally in users’ web browsers, supporting a large number of concurrent users. Source code and an interactive online demo are available at https://github.com/Visual-Intelligence-UMN/GNN-101.
2. Related Studies
2.1. Visualizations for Learning Algorithms and Machine Learning Models
Interactive visualization has a long history of being used for understanding complicated algorithms and machine learning models. Using visualization for algorithms can be traced back to the 1970s (Hundhausen et al., 2002). By visually representing the dynamic nature of computations (e.g., sorting, searching), visualization facilitates the exploration of algorithmic behavior and enhances the understanding (Stasko, 1990; Gurka and Citrin, 1996). However, these studies focused on more traditional algorithms and may not be directly applied to visualize the complicated mechanism of modern machine learning models.
With the emerging of ML, interactive visualizations of ML are gaining wide popularity both in academia and beyond. Visualization techniques have been developed for a variety of deep learning models, including CNNs (Liu et al., 2016; Wang et al., 2020), RNNs (Strobelt et al., 2017; Ming et al., 2017), GANs (Wang et al., 2018), and transformers (Yeh et al., 2023; Bycroft, 2024). However, these techniques are specifically designed for models that work with Euclidean data (e.g., images and text) and cannot be directly applied to GNNs.
While recent research has introduced visualizations for GNNs (Jin et al., 2022; Liu et al., 2022), these tools are primarily aimed at AI developers for model debugging and provide limited support for beginners to learn GNNs. This study aims to fill this gap via a GNN visualization tool tailored to the education of beginners.
2.2. Educational Tools for GNNs
The growing application of GNNs has led to the creation of various educational resources for GNNs, ranging from online articles and lecture videos to computational notebooks. These resources generally aim to convey high-level concepts of GNNs through abstract diagrams and mathematical equations. While helpful, their passive nature and lack of interactivity often make it difficult for beginners to fully grasp the complexities of GNNs.
Most relevant to our study are the interactive Distill articles by Daigavane et al. (Daigavane et al., 2021) and Sanchez-Lengeling et al. (Sanchez-Lengeling et al., 2021). These articles introduce interactive GNN visualizations but tend to use simplified data and visualizations. For example, the GNN Playground proposed by Sanchez-Lengeling shows the embedding of graphs rather than the complete layer-by-layer computation within a GNN. Daigavane et al.  only illustrated a single layer and utilized scalar numbers to represent node features, which are typically high-dimensional. This gap underscores the need for more comprehensive and realistic visual educational tools.
3. Designing GNN101
Target Users and Usage Scenarios: GNN101 is designed for AI beginners (e.g., students in data science or machine learning) who have a basic understanding of machine learning and are interested in learning about GNNs. GNN101aims to provide an introduction of the foundational concepts of GNNs and complements, rather than replaces, existing educational resources by offering comprehensive and intuitive visualizations of GNNs’ inner workings.
Reviewing Existing GNN Tutorials: To guide the design of GNN101, we analyzed 17 GNN tutorials collected from online searches, including highly-cited distill articles (Sanchez-Lengeling et al., 2021; Daigavane et al., 2021), high-impact GNN courses (Leskovec, 2024; Antonio Longa, 2024; of Amsterdam, 2024), official tutorials from widely-used GNN libraries (Tensorflow, 2024; Geometric, 2024), and highly-rated YouTube videos. We independently encoded the content of these tutorials to summarize which concepts were taught and how were they taught, as shown in Table 1. For the full list of tutorials and their encodings, please refer to the supplementary material.
The existing GNN education resources focus on explaining three main concepts: the input data of a GNN model, the computation process within a GNN process, and the GNN architectures. The three concepts have been covered by all the 17 tutorials we surveyed. For data in GNNs, all tutorials explain graph structure, with 15/17 covering high-dimensional node features. Regarding layer computation, all tutorials explain the inputs and outputs of a GNN layer. Other important concepts include neighboring node aggregation (14/17), neighboring node weights (13/17), and neighboring node sampling (9/17). For overall GNN architecture, tutorials mainly focus on various GNN variants (e.g., GAT, GraphSAGE) (13/17), GNN applications for different tasks (14/17), and the use of non-GNN layers (e.g., MLP) within GNN models (11/7).
Regarding the format of explaining these concepts, most tutorials use a combination of mathematical formulas and abstract diagrams to explain these concepts. Given Python’s prominence in GNN model support (e.g., PyTorch Geometric and DGL), many tutorials (10/17) include code blocks to demonstrate how different layers constitute a GNN model in Python. Surprisingly, only four surveyed tutorials included visualizations of real data. These visualizations primarily depict dimension reduction results of high-dimensional node or graph features. Notably absent are visualizations illustrating the inner mechanisms of GNNs
Design Goals: GNN101 aims to address the limitations of existing tutorials while ensuring the same coverage of the foundational GNN concepts. The design goals were initially developed based on our review and then iterated and refined with four GNN experts throughout the development process. The four GNN experts comprised two professors (E1, E2) who have taught courses on GNNs for more than four years and two GNN researchers (E3, E4) who are main contributors to popular GNN libraries and have extensive experience in educating the user community. E1 and E3 are also co-authors of this paper.
Which Concepts are Taught
Data: 17/17 Inside a GNN Layer: 17/17 Model Architecture: 17/17
Graph
Structure: 17/17
High-Dim
Features: 15/17
Layout Input and Output: 17/17 Aggregation of Neighbors: 14/17 Weights of Neighbors: 13/17 Sampling of Neighbors: 9/17 Non-GNN Layer: 11/17 Different Tasks: 14/17 GNN
Variants: 13/17
How are They Taught
Mathematical Formula: 15/17 Abstract Diagram: 14/17 Python Code: 10/17 Visualization of Real Data: 4/17
Table 1. Reviewing Existing GNN Educational Resources
  1. G.1
    Demystify mathematical formulas: As with most AI models, learning GNNs involves the understanding of various complex mathematical functions involved in the model (e.g., Graph Attention (Veličković et al., 2018), Graph Convolution (Kipf and Welling, 2017), GraphSage (Hamilton et al., 2017)). While most of the surveyed tutorials (15/17) made significant efforts to explain these formulas through diagrams and text annotations, they mainly explained the mathematical formulas but fell short of connecting them to the actual data changes within a GNN. To address this gap, GNN101 proposes interactively linking different parts of a mathematical formula to the corresponding GNN visualizations (subsection 4.3). In addition, users can explore different graph tasks and GNN models and compare the underlying mathematics (subsection 4.4).
  2. G.2
    Integrate Diverse Computations: A comprehensive understanding of GNNs encompasses a wide range of computations, as shown in Table 1. Although these computations are covered by the surveyed tutorials, they are often explained in separate sections, making it difficult to build connections among them. Therefore, GNN101 presents hierarchical levels of detail (subsection 4.1) that smoothly integrate the model architecture, individual layers, and the detailed data transformations. Additionally, complementary views (subsection 4.2) are provided to ensure suitable visualizations for different concepts.
  3. G.3
    Connect abstract demonstration with large scale real data: Direct exploration of real-world examples is crucial for understanding complex machine learning models (Wang et al., 2020; Bycroft, 2024). However, most existing tutorials (14/17) rely on simplified diagrams (e.g., an illustration of a 4-dimensional vector) for demonstrations. While some tutorials (10/17) use python code to access real data and provide high-level data summaries (e.g., node feature dimensions), only 4 incorporate visualizations of real input and output data, leaving no visualizations of the internal workings of GNNs unclear. To respond, GNN101 enables users to explore GNN visualizations with different real datasets (subsection 4.4). It also facilitates the examination of large-scale data through hierarchical levels of detail (subsection 4.1).
4. Visualization Interface
Designed for beginners, GNN101 explains the mathematical operations essential for understanding common GNN variants and tasks. Specifically, it visualizes the forward propagation of a trained GNN model, illustrating how an input graph is transformed for node-, edge-, or graph-level predictions. As depicted in Figure 2, the interface of GNN101 consists of a control panel for selecting input graphs and GNN models across different architectures and tasks (A), an inner model visualization that displays the GNN model’s inner workings (B), and a text panel with an on-boarding tour (C) that guides users in interpreting and interacting with the visualization. In this section, we introduce the main components of GNN101 and explain how they achieve the design goals outlined in section 3.
Figure 2. The interface of GNN101: (A) a control panel for selecting GNN models, tasks, and dataset; (B) an inner model visualization that displays the GNN model’s inner workings; (C) a text panel that guides users in interacting with the visualization (C). The visualization provides hierarchical levels of detail (D1-D4).
4.1. Hierarchical Levels of Details
GNN101 seamlessly integrates hierarchical levels of detail, from the overall model architecture to individual layers and the intricate data transformations within each layer.
First, users can observe the architecture overview next to the GNN model selector, as shown in Figure 2.D1. Different GNN tasks (e.g., node or graph classification) and GNN variants (e.g., graph convolution or graph attention) will lead to different model architectures.
Second, users can view a more detailed visualization that shows the outputs of each layer (Figure 2.D2). Clicking on a layer in the architecture overview highlights the corresponding layer in the detailed visualization. Users can also observe the connections between layers and how node features (i.e., a high-dimensional numerical vector) are updated by learning from the features of their neighbors layer by layer.
Third, when users click on a specific layer, GNN101 visualizes the detailed computation within that layer in the style of a horizontal flow chart, as shown in Figure 2.D3 and Figure 3.b. Inputs, outputs, and learnable parameters of this layer are visualized as heatmaps (Figure 3.B1), where each cell’s color represents its value. Curves connecting these heatmaps illustrate the computation process. Internal results are also displayed as heatmaps to break down a complex computation process into multiple steps, making it easier to interpret. Icons on the curves indicate specific types of computations (Figure 3.B2), while curves without icons represent the addition of factors. The thickness and color of these curves encode the corresponding multiplication factors.
Lastly, to gain a deeper understanding of each step of the computational process within a layer, users can hover over cells in the heatmaps, which represent the inputs and outputs of the break-down computations. A pop-up window will then reveal the specific calculations used to determine the value of the selected cell, as shown in Figure 2.D4.
The hierarchical levels of detail in GNN101 allow users to interpret GNN concepts with different levels of granularity (G.2). Additionally, by revealing data and computation processes hierarchically, GNN101 enables effective visualization of large-scale real-world data without overwhelming users (G.3).
Figure 3. Matrix View: The matrix view complements the node-link view (Figure 2) and provides similar click-to-expand interactions (a-b). The computation process inside a layer is visualized as a horizontal flowchart, where heatmaps represent vectors and matrices (B1), and the connecting curves illustrate the computation process (B2).
4.2. Complementary Views
GNN101 provides both a node-link view (Figure 2) and a matrix view (Figure 3) for visualizing GNN’s inner workings.
In the node-link view, the input graph and the intermediate layer outputs are visualized as node-link diagrams, as shown in Figure 2.B. This view not only offers an intuitive representation of the graph structure, but also illustrates the key concepts of a GNN model: updating the features of a graph layer by layer. Users can hover over a node to examine its feature at the corresponding layer. Each feature is visualized as a heatmap, with the color of each rectangle indicating the dimension’s value. At a certain layer, a node’s feature is computed based on its feature and those of its neighboring nodes from the previous layer. Such connections between the selected node and relevant nodes from previous layers are also highlighted when hovering.
In the matrix view, the input graph and the intermediate layer outputs are visualized as matrices, reflecting the exact format in which they are processed within a GNN model. The graph structure is visualized as an adjacency matrix, where each row and column represent a node, and each cell indicates whether the corresponding nodes are connected. Node features are visualized similarly to those in the node-link diagram (i.e., heatmaps), but are arranged horizontally to align with the corresponding rows in the adjacency matrix. Unlike the node-link view, where node features are displayed only upon hovering, the matrix view presents all node features for each layer simultaneously to offer a comprehensive overview (G.2). Users can hover on a node feature to highlight its connections with the relevant node features from the previous layer, mirroring the interaction in the node-link view.
In both the node-link view and the matrix view, users can further investigate the detailed computation process within a layer by clicking to expand the layer, as shown in Figure 2.D2 and Figure 3.b. Non-selected layers will fade out, allowing users to focus on the selected layer while maintaining an overview of the entire model. This click-to-expand feature enables users to examine the overall model architecture and different layers in an integrated and cohesive manner (G.2).
The two views complement each other: While the node-link view offers node-link diagrams for an intuitive representation (Ren et al., 2019), the matrix view provides a comprehensive overview of features and shows data in the exact format used in GNN models. Meanwhile, GNN101 facilitates smooth transitions between two views via consistent interactions (i.e., hover over and click) and visual encoding (i.e., color encoding of the node features).
Figure 4. Bidirectional Math-Visualization Linking: Users can hover over parts of the mathematical formulas to highlight the corresponding visualizations (a), or hover over visualizations to reveal the computation process for obtaining the exact value (b).
4.3. Interactive Math-Visualization Linking
GNN101 offers interactive bidirectional linkage between mathematical formulas and their corresponding GNN visualizations, aiding in the interpretation of complex mathematical concepts (G.1). When users click to expand a layer, the corresponding mathematical formula appears above the visualization. Users can hover over mathematical notations in the formula to highlight the corresponding parts in the visualizations. For example, as shown in Figure 4.a, hovering over the $b$ symbol in the formula highlights the bias vector in the flow chart. Conversely, hovering over visualizations, such as cells in heatmaps, reveals detailed mathematical calculations of exact values, as shown in Figure 4.b. The calculation using actual values provides concrete examples for a better understanding of abstract mathematical formulas. Since this hover interaction only explains the calculation of a single dimension of a high-dimensional node feature, GNN101 also provides an animation that demonstrates the step-by-step calculation for all dimensions.
4.4. Data, Task, and Model Exploration
In the control panel, users can easily explore different GNN models for various tasks and input graphs (G.3). GNN101 provides consistent encoding and smooth transitions to facilitate seamless exploration.
Data: GNN101 provides three graph datasets, a set of chemical compound graphs (Debnath et al., 1991), a social network of a Karate club (Zachary, 1977), and a social network of Twitch players (Rozemberczki and Sarkar, 2021). The graph sizes range from small (e.g., five nodes) to large (e.g., hundreds of thousands of nodes). Graphs are displayed either as a node-link diagram in the node-link view or as an adjacency matrix in the matrix view. For large graphs, GNN101 visualizes only the relevant subgraph associated with a specific prediction, ensuring essential information is conveyed without overwhelming the user with the entire graph. For instance, when predicting a selected edge using a model with two GNN layers, only the two-hop neighbors of the two nodes connected by this edge will be involved in the model inference and visualized accordingly.
Additional Layers for Different Tasks: GNN101 covers common GNN tasks, including node-, edge-, and graph-level predictions. Task-specific layers (e.g., global pooling for graph classification) are also expandable and their inside details are visualized in the similar flow chart style as the GNN layers.
GNN Variants: GNN101 covers three widely-used GNN variants, Graph Attention Networks (GAT) (Veličković et al., 2018), Graph Convolution Networks (GCN) (Kipf and Welling, 2017), and GraphSage (Hamilton et al., 2017), based on the 17 surveyed tutorials. GNN101used a consistent flow chart encoding to visualize these GNN variants while highlighting their distinct designs.
4.5. Implementation
GNN101 is a web-based, open-source visualization tool that can be easily accessed via a web browser without any installation. GNN101 includes pre-trained GNN models and interactive visualizations of these models. The GNN models are pre-trained in Python using Pytorch Geometric (Fey and Lenssen, 2019), covering various graph datasets, tasks, and GNN variants. These pre-trained GNN models are then exported into ONNX format. The visualization is implemented in TypeScript using Next.js, React, and D3.js (Bostock et al., 2011). The pre-trained GNN models are loaded and run in real-time in the user’s web browser using the ONNX Web Runtime (Microsoft, 2024). An anonymized version of the source code is available at https://anonymous.4open.science/r/GNN-101, and an anonymous web demo can be accessed at https://gnn101.vercel.app/.
5. Usage Scenarios
GNN101 is currently being used in three courses that involve the teaching of GNNs at three different universities  [^1] Details omitted for anonymous review. In this section, we show how GNN101 can help a computer science student, Alice, learn various concepts in GNNs based on the reported usage from students.
5.1. Understanding Message Passing
Alice wants to gain a deeper understanding of message passing, the fundamental mechanism underpinning most GNNs (Veličković et al., 2018; Kipf and Welling, 2017; Sanchez-Lengeling et al., 2021). However, the lecture slides and online GNN educational materials primarily explain this concept using diagrams and mathematical formulas, as shown in Figure 1.a, which Alice finds less intuitive and effective. To enhance her learning, Alice turns to GNN101 for assistance.
She begins by exploring the outputs of each layer, which are visualized as a node-link graph, with the features of each node represented as a heatmap. When Alice hovers over a node in the GCNConv2 layer output, GNN101 highlights the selected node, along with its connections to itself and neighboring nodes in the previous layer’s (GCNConv1) output. This interaction helps Alice intuitively grasp the high-level concept of message passing: updating a node’s features by aggregating information from both its neighbors and itself (Figure 2.a).
Curious about the detailed computations of message passing in GCN, Alice clicks on the node to expand the layer and reveal the calculation process as a horizontal flowchart (Figure 2.b). A mathematical formula of the calculation $x^{\prime}_{i}=\sigma(W\sum_{j\in N(i)\cup i}\frac{1}{\sqrt{d_{i}d_{j}}}x_{j}+b)$ also appears to facilitate the understanding. $i$ is the node index, $N(i)$ represents the neighboring nodes of $x_{i}$, $x^{\prime}_{i}$ indicates the updated features for node $i$. On the left of the flowchart, GNN101 shows the relevant node features (heatmaps) from the previous layer, corresponding to $x_{j},j\in N(i)\cup i$ in the formula. The aggregation of these node features $\sum_{j\in N(i)\cup i}\frac{1}{\sqrt{d_{i}d_{j}}}x_{j}$ is visually represented by curves connecting each feature to the aggregation result. The color and thickness of the curves illustrate the weighting factor $\frac{1}{\sqrt{d_{i}d_{j}}}$. The summation result is then multiplied by a weight matrix $W$, with an icon indicating the matrix multiplication, and produces an updated feature. This updated feature is then added with a bias vector $b$ and pass a non-linear activation function $ReLU$, generating the output feature $x_{i}=\sigma(W\sum_{j\in N(i)\cup i}\frac{1}{\sqrt{d_{i}d_{j}}}x_{j}+b)$.
Lastly, to refresh her memory on matrix multiplication and the ReLU activation function, Alice hovers over the results of these calculations, triggering tooltips that explain the process and demonstrate how specific values are computed (Figure 2.D4).
Figure 5. Comparing GNN variants: GNN101 supports the comparisons between GCN (a), GAT (b), and GraphSAGE (c) by visualizing how each model aggregates information from neighboring nodes. Curved edges indicate that parts of the visualization have been omitted from the figure due to space constraints.
5.2. Comparing GNN Variants
After understanding message passing and its implementation in GCN, Alice recalls that message passing is implemented differently across various GNN variants. The lecture slides offer only mathematical formulas for direct comparison of these GNN variants, which limits her ability to fully grasp the differences. Therefore, Alice turns to GNN101to examine other two GNN variants: GAT and GraphSAGE.
Alice starts by selecting GAT in the control panel and clicking on a layer to view its internal computations. The computations are visualized using a horizontal flowchart similar to the one for GCN but with differences in how the input features are connected and aggregated. Unlike GCN, which uses the node degree $\frac{1}{\sqrt{d_{i}d_{j}}}$ as an aggregation factor, GAT computes an attention score $\alpha_{i,j}$, which is visualized with gradient colors for distinction and interactivity. As shown in Figure 5.B2, Alice clicks on an attention score, revealing a step-by-step breakdown of the computation through three mathematical formulas. To aid comprehension, GNN101 includes text annotations and heatmap visualizations to indicate the learnable weight matrix $W$ and the attention $a$ for this layer.
Next, Alice selects GraphSAGE in the control panel and clicks on a layer to view its computations. Sampling icons (Figure 5.C1) appear on the curves connecting and aggregating input features, highlighting the neighbor sampling process, a key innovation of GraphSAGE. Clicking on these icons triggers tooltips that provide textual explanations of the sampling method used in GraphSAGE.
Alice concludes that the primary differences among the three GNN variants lie in how they select and aggregate information from neighbors, which are visually represented by the curves connecting input features to intermediate summation results (Figure 5.a, B1, and c).
Figure 6. Learning different GNN tasks: The final one or two layers handle task-specific processing for graph classification (a), node classification (b), and edge prediction (c). These additional layers also support click-to-expand interactions, similar to the GNN layers. Curved edges indicate that parts of the visualization have been omitted from the figure due to space constraints.
5.3. Learning Different GNN Tasks
After understanding message passing and its implementation across GNN variants, Alice is eager to explore how GNNs solve various tasks in real data. In GNN101, Alice uses the control panel to switch between three common GNN tasks: graph classification, link prediction, and node classification. The matrix view enables Alice to examine the outputs of all layers at the same time (Figure 3.a). She notices that the first several layers of GNNs for these tasks are very similar. Each layer updates the graph node features by learning from its neighbors.
Alice observes the main differences occur in the final one or two layers for task-specific processing (Figure 6). In graph classification (a), a global pooling layer aggregates all node features of the graph and outputs a single feature, which is used to generate the graph classification via an MLP layer. Alice hovers over this layer to see the detailed computation: an element-wise average across dimensions (Figure 4.b). In node classification (b), an MLP layer is applied to each node feature to generate predictions. In edge prediction (c), the features of the two nodes involved in the predicted edge are fed into a prediction layer that outputs a prediction score. Alice clicks on the prediction layer and sees a dot product of the two node features used (C1).
Alice also observes the representation power of GNNs by directly modifying the input data and seeing how a GNN updates the node features layer by layer to generate an accurate prediction.
6. User Interviews
Evaluating educational visualization tools presents challenges due to the long-term nature of learning complex machine learning concepts. Additionally, comparing GNN101 to baselines is difficult because GNN educational visualizations are largely absent. The closest examples are two interactive articles (Sanchez-Lengeling et al., 2021; Daigavane et al., 2021), both of which are lengthy and require over 40 minutes to read even without interacting with their visualizations.  [^2] Based on an average reading speed of 238 words per minute (Brysbaert, 2019). Therefore, we followed practices from previous studies on educational visualizations (Wang et al., 2019, 2020) and conducted an observational study and interviews to assess the usability of GNN101 and to gain insights into user interactions with it.
6.1. Participants
We recruited 12 participants (10 males, 2 females, average age 23.8) through snowball sampling. Given that learning GNNs from scratch is a long-term process, often requiring more time than a typical in-lab interview (30–60 minutes), our study focused on two types of participants: 1) Teaching assistants (TAs) for GNN-related courses who had experience explaining GNN concepts to students in a conversational, interactive format (e.g., office hours, labs, or tutorials). We have 7 participants in this category, denoted as TA1-TA7; 2) Students who had backgrounds in machine learning (average self-reported skill level of 4.5 on a 7-point Likert scale, where 1 represents Novice and 7 represents Expert) but had not formally learned GNNs. We had 5 participants in this category, denoted as S1-S5. None of the participants had prior knowledge of this project.
6.2. Procedure
We conducted the study with participants one-on-one via Zoom. Before the study, each participant completed an intake form that included demographic questions and optional comments on learning GNNs. They also signed a consent form permitting the recording of their audio and screen activity. Each study began with a 5-minute introduction to GNN101 and an overview of its key functions. After the tutorial, participants freely explored GNN101 in their web browsers, following a think-aloud protocol. A manual outlining the main functions of GNN101 was available for reference, and participants could ask questions at any time during the process. Each session concluded with a questionnaire (same as the one used in CNNExplainer (Wang et al., 2020)) and a semi-structured interview. The sessions lasted 30 to 50 minutes, and each participant was compensated with a \$10 Amazon gift card.
6.3. Observations
Figure 7. Results of the post-study questionnaires: (A) User ratings on system usability; (B) User ratings on various components. Ratings were based on a 7-point Likert scale, where 1 indicates strongly disagree and 7 indicates strongly agree.
Overall, participants stated that GNN101 addresses a strong demand for visualization when learning GNNs. For example, TA6 stated “unlike images that look like grids, the unique structure of graph makes it hard to imagine how the data flows through the model… definitely a very useful visualization”. TA5 emphasized the importance of visual aids in GNN learning, stating, “understanding the core message passing steps of GNNs is pretty visual. I used to reference the animation in this distill.pub article (Sanchez-Lengeling et al. (Sanchez-Lengeling et al., 2021)) a lot when explaining things (to students)”. S2 reported “You guys are more interactive, more comprehensive (compared with (Sanchez-Lengeling et al., 2021))…I think it’s pretty helpful, for sure”. TA7 commented “I think this is the best tool out there for learning graph neural network”.
Our observation also showed that GNN101 provides a user-friendly, easy-to-use interface. After the short tutorial, most of the participants (9/12) were able to conduct the free exploration without asking the interviewer questions or referring to the manual. Two participants sought clarification on the visual encoding to confirm their interpretation of the adjacency matrix, which was correct. Even though some participants initially expressed confusion about certain interactions (e.g., saying “Wait, what does this mean?” or repeatedly hovering over specific areas), they were able to resolve these issues independently. For example, after asking a question, TA5 immediately followed by “Oh, I see now, it means the neighbor sampling. That is nice”, before the interviewer had a chance to intervene.
This overall positive attitude is also reflected in the post-study questionnaires, as shown in Figure 7. On a 7-point Likert scale, all average ratings are above or equal to six. Notably, three participants mentioned that they would be continuing as TAs in the upcoming fall semester and asked permission to use GNN101 for the students.
User Interactions: We observed some consistent patterns in how participants interacted with GNN101.
First, most participants (9/12) followed an overview-to-detail approach for exploration: they began by exploring the overall model structure, then selected specific nodes to observe the message passing, and finally hovered over individual components to examine the underlying mathematical operations. Such an interaction flow demonstrates the importance of the hierarchical level of details, as noted by TA4, “I like it … if we present all of the information at once, it might be confusing to know what to focus on first”.
Second, we also observed a tendency to move from intuitive to concrete. All participants chose the node-link view to start their exploration, likely due to its default status and the intuitive representation of graph data. As TA4 commented, “that’s (message passing on graph view) super clear… when I worked as a TA, I found that is where other students have trouble understanding.” Participants then switch to the matrix view to “see the outputs of all layers” and “connect theoretical concept with practical GNN data flows” (TA7).
Third, some participants initially struggled with visualizations but found clarity through the integrated math-visualization linking. For instance, S4 initially felt confused when viewing the message-passing visualization. However, S4 understood the visualization after interacting with the formula, which describes the message passing as the aggregation of neighboring node features. S4 remarked: “This is probably my favorite feature.”
Comparing with Existing Tools: In the post-study interviews, participants frequently compared GNN101 with other GNN educational resources, including interactive articles on distill.pub (Sanchez-Lengeling et al., 2021; Daigavane et al., 2021), lecture slides (Leskovec, 2024), whiteboard drawing, lecture videos, and YouTube educational videos. It’s important to clarify that GNN101is not intended as a complete replacement. For instance, it doesn’t provide the comprehensive textual introductions found in (Sanchez-Lengeling et al., 2021; Daigavane et al., 2021) or the programming implementation details offered by Pytorch Geometric tutorials (Antonio Longa, 2024) (as mentioned by TA5). At the same time, most participants found GNN101 presents a unique advantage, especially through the hierarchical levels of detail and the integration of mathematical formulas with visualization. S3 remarked: “Those are the two things (the hierarchical levels of detail and complementary views) that make it more useful than a YouTube video or the Distill article.”
Suggestions: The most common suggestion we received was to incorporate additional features that support more advanced concepts in GNNs. These suggestions included inductive and transductive learning, the aggregation of both node and edge features, and comparisons among GNNs and other neural network architectures (e.g., MLPs and CNNs). Another frequent suggestion was to allow users to interactively adjust GNN parameters, architecture, or input data and observe the resulting changes. Other suggestions were mainly about minor style changes such as increasing the font size, making the control panel resizable, and adding model training information. These suggestions reflect the participants’ appreciation of the already-implemented functionality and their desire to extend GNN101 to encompass broader concepts. Importantly, even while proposing these enhancements, participants who offered such suggestions agreed that the current version of GNN101 already provided comprehensive support for foundational GNN concepts.
7. Discussion
7.1. Design Implications
An important design implication emerging from this study is the value of animated visualization transitions in educational contexts. Based on the design iterations with experts, GNN101includes a variety of animated transitions, including visual transportation of node features from the node-link graph to the flowchart, step-by-step animations of matrix multiplication, and progressive reveal of the computation flowchart from left to right. Even when these animations do not necessarily convey additional information, they offer an engaging and attractive learning experience and help users connect various visual components across different levels of detail. While previous studies have shown the effectiveness of animated transitions in learning visualization (Ruchikachorn and Mueller, 2015; Heer and Robertson, 2007), their applications are rarely discussed in the emerging VIS4ML community. This discrepancy may reflect the distinct visualization requirements between educational and analytical contexts. As T7 noted, “it (the animated transition of node features) can be annoying if I am analyzing a GNN model for my own research, but a big plus for explaining things to students”.
Another crucial implication of our study is the integration of mathematical formulas into AI model visualizations. Surprisingly, this seemingly straightforward linkage is often absent in current AI visualizations. Traditionally, visualizations have been proposed as alternatives to the complex mathematics of AI, offering a different perspective on model functionality. However, without explicit connections between visualizations and their underlying mathematical foundations, users, especially new beginners, may struggle to bridge the gap between textbook formulas and visual representations. While a few examples of math-visualization linkage exist, such as CNNExplainer (Wang et al., 2020) demonstrating specific value computations and Daigavane et al. (Daigavane et al., 2021) using color-coding to link node features with their mathematical notations, these implementations are simple and often limited in scope. GNN101 introduces bidirectional math-visualization linking. This approach not only aids in demystifying AI models but also strengthens the connection between theory and practice.
7.2. Limitations and Future Work
Our current implementation of GNN101 has shown promise in enhancing GNN education, but there are several areas for improvement.
First, the current version only supports a list of provided graph data. While users can upload a graph as a JSON file to be predicted and visualized, it is hard for users to update graph data in the same format supported in GNN101. Future iterations of GNN101 could incorporate dynamic graph modification capabilities, allowing users to modify node and edge in real-time and observe immediate GNN responses to these modifications.
Second, while the current evaluation demonstrates the usability and effectiveness of GNN101, the evaluation is limited to interviews with a limited number of participants. To better assess the educational impact of GNN101, we are planning on a large-scale deployment of GNN101 in three GNN courses across three universities, which will enable long-term observation and collaboration of diverse data. Such an extensive evaluation can provide valuable insights for future studies in visualization for AI education.
Third, even though the current version of GNN101 provides comprehensive coverage of the foundational concepts in GNNs, we received many suggestions for additional advanced features in the interviews, reflecting the complexity of GNNs. In the future, we plan to evolve GNN101 into a modular framework and a collaborative platform that will allow the community to contribute visualizations for a wider variety of GNN models, tasks, and datasets.
8. Conclusion
This study proposes GNN101, an interactive visualization tool for learning GNNs. Through a review of existing GNN educational materials and close collaboration with experts in teaching GNNs, we designed and implemented four main features to facilitate the learning process: hierarchical level of details, directional math-visualization linking, complementary views, and exploration of data, models, and tasks. We currently deploy GNN101 in three GNN-related courses across three universities. The results from this deployment, including reported usage scenarios from students, interviews with TAs, and advanced ML learners, show the usability and effectiveness of GNN101. We believe that the proposed visualization techniques and the derived design lessons have implications beyond GNNs, inspiring future studies on visualization tools for AI education.